OpenStudy (gio333):

When a product is raised to a power, it is equal to the product of each factor raised to that power. Show that this rule does not apply for a sum or difference to a power using the problem (5 – 3)^2. You may indicate an exponent in your answer with ^. For example, 3x2 as 3x^2.

OpenStudy (anonymous):

(5 - 3)^2 = (2)^2 = 4 (5 - 3)^2 != (5)^2 - (3)^2 = 25 - 9 = 14

OpenStudy (anonymous):

Where != means is not equal to.

OpenStudy (anonymous):

(5 - 3)^2 = (5 - 3) * (5 - 3) = (2) * (2) = 4

OpenStudy (anonymous):

Make sense?

OpenStudy (gio333):

yeah it does

OpenStudy (anonymous):

Cool. Take care and don't forget to mark a Best Response. Thanks!